briannamilcic:

can some help on this question The table below shows the data for a random sample of fish that were collected from and later released into a lake: Type of Fish Catfish Grass Carp Bluegill Number of Fish 52 61 87 How many grass carps are estimated to be present if a sample of 800 fish is collected from the lake?

Florisalreadytaken:

Type of Fish Catfish Grass Carp Bluegill Number of Fish 52 \(\color{lightskyblue}{61}\) 87 that is a total of \( 52 +61+ 87 = \color{tomato}{200} \) fish "How many grass carps are estimated to be present if a sample of 800 fish is collected from the lake? " we can set up a proportion: \[ \frac{\color{lightskyblue}{61}}{ \color{tomato}{200}} =\frac{x}{800} \] \[ \ \ \ \ \ \ \ \Updownarrow \] \[ 61\times 800=200 \times x \] \[ \frac{61\times 800}{200}=x \Rightarrow \frac{61\times \cancel{800}}{\cancel{200}}=x \] \[ \frac{61\times 4}{1}=x \] \[ \frac{244}{1}=x \Rightarrow \frac{244}{\cancel1}=x \] \[ 244=x \] that would be the answer.

ItzAgar:

Do you still need help on this question?